OpenStudy (anonymous):

3, 10, 29, 74, ? find the next term?

OpenStudy (anonymous):

option are (a) 125 (b) 173 (c) 111 (d) 184

OpenStudy (anonymous):

along with explanation plz?

OpenStudy (anonymous):

A_(n-1)*3 - 1 = An

OpenStudy (anonymous):

\[A_{n-1}*3 -1\]

OpenStudy (anonymous):

i didnt get it...will u plz show by putting numerical value.

OpenStudy (anonymous):

Disregard above answers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you recheck your series

OpenStudy (anonymous):

*sequence

OpenStudy (anonymous):

bingo i got an answer myself.

OpenStudy (anonymous):

its 173.

OpenStudy (anonymous):

3*2+4=10 10*2+9=29 29*2+16=74 74*2+25=173.

OpenStudy (anonymous):

thanks for ur time guys/